OpenStudy (pugtato):

I understand what direct variation is but now I need help solving a different question and im confused again... Which equations show a direct variation between x and y when solved for y? Choose all answers that are correct. A. y-9=2x+9 B. y=4/7x C. 4+y=1/3x+4 D. y=3/x

OpenStudy (pugtato):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let's focus on y-9=2x+9 for now

jimthompson5910 (jim_thompson5910):

how would you isolate y?

OpenStudy (pugtato):

+9 and cancel it out?

jimthompson5910 (jim_thompson5910):

yes add 9 to both sides

jimthompson5910 (jim_thompson5910):

y-9=2x+9 y-9+9=2x+9+9 y = 2x + 18

OpenStudy (pugtato):

so y=2x+18

OpenStudy (pugtato):

oh you already put it lol

jimthompson5910 (jim_thompson5910):

that's ok. So is y = 2x+18 in the form y = k*x ?

OpenStudy (pugtato):

I think if there's no 18, then yes

jimthompson5910 (jim_thompson5910):

so because of that added 18, it means it's not in direct variation form

OpenStudy (pugtato):

So A can't be one of the answers

jimthompson5910 (jim_thompson5910):

agreed

jimthompson5910 (jim_thompson5910):

if choice B is \[\Large y = \frac{4}{7}x\] then is it in y = k*x form?

OpenStudy (pugtato):

Yes

jimthompson5910 (jim_thompson5910):

good, so k = 4/7 in this case

OpenStudy (pugtato):

And for C, i'll try to do the same thing we did for A.

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (pugtato):

Cancel out the 4 on both sides and you get y=1/3x which is in the form of y=k*x, right?

jimthompson5910 (jim_thompson5910):

yep k = 1/3 here

OpenStudy (pugtato):

So what about D?

jimthompson5910 (jim_thompson5910):

D is in the form y = k/x where k = 3 in this case. This is inverse variation (not direct variation)

OpenStudy (pugtato):

Oh like the other one we did earlier, so D isn't an option... so the answers are A and C

OpenStudy (pugtato):

*B and C

jimthompson5910 (jim_thompson5910):

yes B and C are the two answers

OpenStudy (pugtato):

so the next question is asking for the constant of variation, how do I find that?

jimthompson5910 (jim_thompson5910):

it's the value being multiplied by the x example: y = 3x has a constant of variation of k = 3

jimthompson5910 (jim_thompson5910):

focus on the column with x = 1 and y = 3 y = k*x 3 = k*1 ... replace x with 1 and y with 3 solve for k

OpenStudy (pugtato):

1?

jimthompson5910 (jim_thompson5910):

3 is equal to SOMETHING times 1 what is that SOMETHING?

OpenStudy (pugtato):

1/3?

jimthompson5910 (jim_thompson5910):

3 = k*1 so k = 3

jimthompson5910 (jim_thompson5910):

k = 3 is the constant of variation

OpenStudy (pugtato):

Ohh It's 3 for the rest of them too if you solve them the same way

jimthompson5910 (jim_thompson5910):

yeah notice how if x = 2 then y = 6. You'd get the same value of k